# GPU-Accelerated Algorithms for Process Mapping thanks: Henning Woydt is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – DFG SCHU 2567/6-1

### v1

@@ V-A Results
Comparison to GPU-based Algorithms. While no other GPU-based algorithms currently exist for solving the GPMP, we can compare our algorithms to Jet [18]. Note that Jet solves the graph partitioning problem, which corresponds to using a distance vector of $D=1:\ldots:1$. Since we test the distance $D=1:10:100$, Jet’s resulting partitions are unfit and the corresponding communication costs are worse than those of our algorithms. In terms of runtime, however, Jet is comparable to GPU-IM as they both follow the multilevel scheme. We base our comparison on the large graph instances ($\geq 1$ million vertices). Using GPU-IM as the baseline for runtime and communication costs, we observe that Jet and Jet-ultra incur additional communication costs of 42.142.1% and 38.838.8% on average respectively. In terms of runtime, Jet-ultra is comparable to GPU-IM, with a geometric mean speedup of $1.03$, whereas Jet achieves an average speedup of $1.33$. We attribute this performance gap primarily to the 4040% matching requirement of our algorithm, compared to the 7575% in Jet. This allows Jet to produce smaller coarse graphs in fewer levels, and thereby reduce runtime. While we could tune our constants for more speed, we observed that it negatively impacted solution quality. Our algorithms GPU-HM and GPU-HM-ultra further improve solution quality but are worse in runtime than GPU-IM (see Figure 1), which translates to this GPU comparison.
VI Conclusion
@@
Process mapping assigns vertices of a task graph to the processing elements (PEs) of a supercomputer to balance workload and minimize communication costs. Motivated by the success of GPU-based graph partitioners, we proposed two GPU-accelerated algorithms for this problem. The first, GPU-HM, applies hierarchical multisection [47] using the GPU-based partitioner Jet [18]. The second, GPU-IM, integrates process mapping directly into a multilevel framework, adapting Jet’s label propagation and rebalancing. GPU-HM-ultra achieves up to $300\times$ speedups and is on average a magnitude faster than the state-of-the-art SharedMap [44], with only 11.511.5% higher communication costs and therefore outperforming IntMap-S [15]. GPU-IM delivers extreme speedups (mean $77.7\times$, max $598.5\times$) at lower solution quality, however it still outperforms the Fast configuration of SharedMap. Future work will focus on improving solution quality without sacrificing performance e.g., enhancing label propagation filters, refining coarsening and initial partitioning, and exploring more efficient GPU-based partitioners.

### v2

@@ 5.4 GPU-based Comparison
While no other GPU-based algorithms currently exist for solving the GPMP, we can compare our algorithms to Jet [19]. Note that Jet solves the graph partitioning problem, which corresponds to using the distance vector $D=1:\ldots:1$. Since we test the distance $D=1:10:100$, Jet’s resulting partitions are unfit and the corresponding communication costs are worse than those of our algorithms. In terms of runtime, however, Jet is comparable to GPU-IM as they both follow the multilevel scheme.
Jet’s determined partitions incur on average $45.3\%$ additional communication cost over those by GPU-IM. Compared to SharedMap-S this is an average increase of $90.3\%$. As explained it is expected that the produced partitions are unfit for process mapping, which highlights the need for a dedicated solver. Note that the quality of the ultra configuration is even worse. This can be explained because decreasing the edge-cut does not necessarily decrease the communication cost. Trading two cut edges for one cut edge is always beneficial (assuming uniform edge weights) when optimizing the edge-cut, however this is not inherently true for process mapping. Trading two cut edges on neighboring PEs for one cut edge that connects two islands has an overall detrimental effect. In terms of runtime, GPU-IM is slightly faster than Jet. On small graphs, with fewer than one million vertices, the geometric mean speedup is $1.43\times$, with a minimum and maximum speedup of $0.21\times$ and $1.95\times$. For larger graphs, with more than one million vertices, the mean speedup increases to $1.56\times$. The minimum and maximum speedup are $1.21\times$ and $2.22\times$. We primarily attribute the small speed gains to the use of the extended CSR format and miscellaneous technical improvements. In summary, GPU-IM has a geometric mean speedup of $1.47\times$ over Jet and therefore delivers equal speed to current state-of-the-art GPU-based graph partitioners.
6 Conclusion
@@
Process mapping assigns vertices of a task graph to the processing elements (PEs) of a supercomputer to balance workload and minimize communication costs. Motivated by the success of GPU-based graph partitioners, we proposed two GPU-accelerated algorithms for this problem. The first, GPU-HM, applies hierarchical multisection [48] using a reimplementation of the GPU-based graph partitioner Jet [19]. The second, GPU-IM, integrates process mapping directly into a multilevel framework, adapting Jet’s label propagation and rebalancing. GPU-HM-ultra achieves up to $934\times$ speedups and is on average $20$ times faster than the state-of-the-art algorithm SharedMap [45], with only $12.0\%$ higher communication costs and thereby outperforming IntMap-S [16]. GPU-IM delivers extreme speedups (mean $1454\times$, max $12376\times$) at lower solution quality, however it is still comparable the Fast configuration of both SharedMap [45] and IntMap [16]. Future work will focus on improving solution quality without sacrificing performance, e.g., enhancing label propagation filters and refining initial partitioning.